ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35,46,76 Morimoto, Hiroko, Meiji University, Japan On the Existence of Steady Navier-Stokes Flow in 2D Tube-like Domain under General Outflow Condition Abstract: We discuss about the existence of steady Navier-Stokes flow in 2 dimensional tube containing some obstacles, symmetric with respect to the axis. We don’t suppose the flux vanishes on the boundary but suppose only general outflow condition. Assymptotic behaviour of the solution is also discussed.